# General relativity and the double-zero eigenvalue

### v1

@@ 3.1 General remarks
There are two main cases to consider, very different qualitatively, for a gravitating system modelled by a dispersive (i.e., structurally unstable) set of dynamical equations:
@@
This is in sharp contrast to the possible behaviours of structurally stable, hyperbolic systems as discussed above. We say that gravitational equations are of infinite codimension in Case 1, and of finite codimension in Case 2. Although one does not rule out the existence of gravity problems belonging to Case 1, in any encountered problem of infinite-codimension one would typically obtain overdetermined problems and following R. Thom [1], chap. 3, one expects to be able to show that there exists a finite order such that ‘almost all’ enumerations at some smaller order become determinate at that order. This means that in this case higher-order terms beyond that order may be neglected, and we may treat this case as one of finite codimension.
@@
The case of finite codimension, say equal to $s\leq 4$, is the main case we shall consider below. In fact, we shall advance a point of view and various results that point to the finite codimension answer, and for various reasons explained below-while we keep an open mind-we also expect that in most cases of interest the codimension for gravity will be small (up to four).
Let us assume for concreteness that we are given a gravitating system governed by the Einstein equations (for example, a Friedmann-Robertson-Walker (FRW) universe with a perfect fluid source-the following argument remains applicable in any modified gravity and effective string theory situation, etc). There are two basic issues on which our approach to gravitational bifurcations focuses: firstly, the question about finite determinacy: to what extent the low-order terms in a Taylor expansion of the gravitational bifurcation problem completely determine its qualitative behaviour regardless of the higher-order terms that may be present? Secondly, the question of unfolding: how does a gravitational bifurcation depend on parameters? The natural area to formulate both questions and search for complete answers is provided by the elaborate mathematical framework of bifurcation theory.
3.2 Objectives
We may now describe the main objectives of the present approach in a somewhat more precise language. In any gravitational context for example, cosmology, black holes, modified gravity, or effective string theory, we aim to a complete realization of the program of bifurcation theory for a variety of gravitational problems. This program is one of accomplishing the following three objectives that offer a complete portrait of all possible qualitatively different behaviours for a given gravitational problem (the question of how to accomplish these objectives is analysed in Section 4, while examples of how these objectives have worked out so far may be found in Section 5).
3.2.1 First objective
Formulation: For typical gravitation problems, demonstrate that they are bifurcation theory problems.
This objective deals with the first main difficulty to realize the bifurcation theory program, namely, to identify the type of degenerate behaviour present in the given gravity problem and separate it from any remaining hyperbolic one.
We have to demonstrate that the given set of gravity equations indeed constitute a bifurcation problem. In this regard, we need to solve the following problems:
@@ 3.2.1 First objective
3.2.2 Second objective
Formulation: Given the results of Objective-1 for a given gravity problem, provide complete answers to the finite determinacy and unfolding questions.
@@ 3.2.2 Second objective
3.2.3 Third objective
Formulation: Develop the physics of versal unfoldings associated with the original gravity problem.
While the first two objectives are almost purely mathematical and lead to the constructions of the topological normal forms and the bifurcation diagrams, in this objective the main question is to discover the physical implications stemming from these constructions as regards the original physical problem.
The importance of this objective naturally leads to new physical effects totally absent in the original system. This clearly follows from the fact that the resulting constructions from the first two objectives are very closely interlocked with, and follow from, the degeneracies of the original system, all of which are by necessity ignored in a hyperbolic treatment.
To motivate this objective, let us start by emphasizing the fact that there are three kinds of parameters in the versal unfolding (and associated bifurcation diagrams):
@@ 3.3.1 Bifurcation approach compared
3.3.2 Gravitational completion
It follows that a bifurcation theory approach to gravitation leads to a field-theoretic approach to gravitating systems that leads to the discovery of novel aspects of gravitation that cannot be found without it, and also provides the means for their further study. It implies possibly observable effects and signatures, definite predictions which would be difficult to capture otherwise but ready to be confirmed or falsified.
This aspect of the present work will also be important in physical cosmology starting from bifurcation results. This is so for the following three main reasons:
@@ 3.3.2 Gravitational completion
  • Novel mathematical problems appear by necessity. Bifurcation theory is in a sense a purely algebraic approach to gravitational problems in which the resulting equations fully determine the solutions. For example:
@@
This is where the importance of a gravitational bifurcation approach lies. This approach represents the first few steps towards a more comprehensive theory of ‘gravitational completion’, and independently, an approach to a problem first posed in Ref. [43]. For it provides a reliable path from the large body of known and well-understood results related to a hyperbolic approach, to the description of a consistent (i.e., structurally stable) framework, a new but yet unknown structure sought-for in this approach, which contains all versal unfoldings of a given original theory. For example, what are the versal unfoldings, and the precise relations between them, of the Einstein-Maxwell, Einstein-Dirac, or Wheeler-DeWitt equations?
4 Dynamics and bifurcations
We shall now describe in some detail some of the main methods of bifurcation theory we shall use subsequently in Section 5. We have developed the introductory material in the few subsections, while later in this Section we discuss additional, somewhat more advanced methodology topics.
Consider a set of dynamical equations that govern the evolution of a gravitating system, we shall call this set the original problem. For example, any of the problems analyzed in Section 5 is well suited.
4.1 Dynamical systems formulation
We assume that there is a change of variables that brings the differential equations that define the original problem in the following dynamical system form,
$\dot{w}=G(w,\lambda),$ (4.1)
Another reason we add this as a separate step belonging to the bifurcation sequence is to emphasize the fact that for our purposes it is important that the new variables $w$ that bring the original system of equations to the form (4.1) are functions of only the ‘phase space’ variables of the original problem, say, $y$, i.e., $w=w(y)$, and not of the distinguished parameters of the problem (when they exist). In other words, we need to ensure that any parameter present in the original problem survives in the form (4.1) and is not absorbed in the transformation.
4.2 Jordan form of the linear part
Let us suppose that (4.1) has a nonhyperbolic equilibrium at $(w,\lambda)=(w_{0},\lambda_{0})$ (in case of more than one nonhyperbolic equilibria, we repeat the procedure for each one of them). We shall assume for simplicity that there are no distinguished parameters (the other case is somewhat more complicated, but the main ideas are qualitatively similar).
$\dot{v}=DH(0)v+\bar{H}(v),$ (4.2)
$\dot{X}=JX+F(X),$ (4.3)
Remark 4.1
Once we have completed the steps in the previous two subsections, we have already gained new and valuable information about the original problem.
Example 4.1
Consider two different gravity problems for which the application of the previous two subsections gives the following two results for the Jordan forms $J_{1},J_{2}$ of their linear parts respectively when we write them in the form of Eq. (4.3):
Which of the two gravity problems with linear part as above is more degenerate? The degree of degeneracy is roughly given by the number of zero eigenvalues in the Jordan forms, the higher this number the greater the degeneracy, and in this case, it is one for the linear part $J_{1}$ and two for $J_{2}$.
4.3 Centre manifold reduction
Then the centre manifold reduction theorem (cf. e.g., [11], chaps. 1, 2) states that: (1) near the (nonhyperbolic) origin there exists an invariant graph (manifold) $y=h(x)$, the centre manifold, such that the dynamics given by Eq. (4.3) is equivalent to the dynamics of the reduced $c$-dimensional system,
$\dot{x}=Ax+f(x,h(x)),$ (4.5)
where $f$ is the $x$-component of the nonlinear part $F(X)$ in Eq. (4.3), that is depending only on the dispersive dimensions $x$; (2) the two solutions, namely, $x(t)$ on the centre manifold and $(x(t),y(t))$ of the system (4.3) are identical up to exponentially small terms, while their stabilities are the same; and (3) the centre manifold can be computed to any finite order of $x$ for small solutions by using the ‘tangency condition’: $\dot{y}-Dh(x)\dot{x}=0$.
$\dot{x}=Ax+f(x,h(x,\lambda),\lambda),\quad\dot{\lambda}=0.$ (4.6)
The centre manifold reduction theorem implies a reduction of the dynamics of Eq. (4.3) down to acquiring the dimensionality of the centre manifold ($c<n$). (We note that the reduction theorem with $\lambda$ goes through similarly because the added $k$-dimensions are along the centre directions and have no dynamics.)
For example, suppose we have a gravitational system which is brought to the form (4.3) after the application of subsections 4.1, 4.2, with the $n$-dimensional Jordan matrix $J$ being of the form of the matrix (4.4a) having one zero and $(n-1)$ eigenvalues equal to $-1$. Then the reduction theorem implies that this problem can be reduced to a one-dimensional system of the form Eqns. (4.5) or (4.6) depending on the existence of distinguished parameters in the original system.
Bifurcation theory assumes that all systems are already reduced to a $c$-dimensional centre manifold by the method of this subsection, and deals with dynamical equations in the form (4.3), or, that in Eq. (4.6) (that is after centre manifold reduction).
4.4 Normal form
In subsection 4.2, we simplified the linear part of the vector field and now we proceed to simplify the nonlinear part $F(X)$ in Eq. (4.3) (in the case of distinguished parameters, we simplify Eq. (4.6) by an analogous but more involved process). Let us assume the matrix $J$ has $n$ distinct eigenvalues $\nu_{1},\dots,\nu_{n}$ (more general cases are treated in a more elaborate but essentially similar process).
We say that the eigenvalue vector $\nu=(\nu_{1},\dots,\nu_{n})$ is resonant if there is a relation of the form $\nu_{s}=m\cdot\nu$, where $m=(m_{1},\dots,m_{n})$, $m_{i}\geq 0,\sum m_{i}\geq 2$, with all $m_{i}$ integers. The number $\sum m_{i}$ is called the order of the resonance, otherwise $\nu=(\nu_{1},\dots,\nu_{n})$ are nonresonant (e.g., $\nu_{1}=2\nu_{2}$ is a resonance of order 2, but $2\nu_{1}=3\nu_{2}$ is not a resonance).
The normal form method of simplification of nonlinear vector fields after centre manifold reduction is essentially due to H. Poincaré who observed that if in Eq. (4.3) the eigenvalues of $J$ are nonresonant, then (4.3) is equivalent to its linear part after a smooth change of the unknown $X$ (this is what happens in the hyperbolic case where we have the best possible conclusion for topological equivalence, i.e., ‘Hartman-Grobman’), and was the fundamental result in Poincaré’s Thesis.
When, however, the eigenvalues of $J$ are resonant, one Taylor expands $F(x)$, and by a sequence of analytic transformations of the form $X=Y+h_{r}(Y),r=2,3,\dots,$ Eq. (4.3) becomes,
where one has succeeded in annihilating all nonresonant terms at each order up to order $r$, and this is the meaning of the word ‘simplification’. This is accomplished by solving a certain linear equation for the $h_{r}(Y)$ (the ‘homological equation’), and at each consecutive order starting with the second, the terms $F_{i}^{\textrm{res}}(Y),i=2,3,\dots$, referred to as resonance terms, represent the remaining unremovable terms after the smooth transformation to the $Y$-coordinates at the given order.
Ideally one wants to remove all inessential (i.e., nonresonant) terms at a given order, and proceed to the next, the final result in Eq. (4.7) contains only the essential terms. The structure of these remaining terms is entirely determined by the linear part of the vector field, while simplifying the terms at order $r$ does not modify any lower-order terms.
One thus obtains an equivalent bifurcation problem to the original, a ‘normal form’ of the field at some desirable order, by constructing at that order a basis of vector fields in the space of homogeneous polynomials which cannot be eliminated. Any term present in the vector field in Eq. (4.3) (or (4.5)) that is not among these unremovable terms of the normal form so constructed can be eliminated.
This ingenious process results in a new vector field which is in many ways much simpler than that given originally and has important applications (for instance in the Kolmogorov-Arnold-Moser theory of quasiperiodicity). For instance, any 2-dimensional vector field that has the matrix $J$ as in (4.4b) as its linear part has normal form $(0,x^{2}\pm xy)^{T}+O(3)$. (This normal form was first discovered by Takens 1974 and Bogdanov in 1975.)
Example 4.2
For the Oppenheimer-Snyder vector field, $J(x,y)^{T}+(0,-3/4y^{2})^{T}$, one can show that the second-order term is nonresonant and so by the above process can be eliminated. Therefore there are no second-order terms in the normal form, and one needs to proceed with a computation of the 3rd-order terms. Setting $X=(x,y))$, the OS vector field becomes,
$\dot{X}=JX+O(3),$ (4.8)
4.5 Versal unfoldings
Although as we showed a centre manifold reduced system can be further simplified, the normal form obtained this way will still be unstable with respect to different perturbations, that is with respect to nearby systems (i.e., vector fields), and so its flow (or, equivalently, that of the original system) will not be fully determined this way.
It is here that bifurcation theory makes its decisive entry, in that it uses the normal form to construct a new system, the (uni-)versal unfolding (or, deformation in other terminology), that is partly based on the normal form but also contains the right number of parameters needed to take into full account all possible perturbations and degeneracies of the normal form system (and so also of the original one).
Consider an equation of the form $g(x,\lambda)=0$, where we assume that $x\in\mathbb{R}^{n},\lambda\in\mathbb{R}^{l}$, for given integers $n,l$, and we may imagine it represents the equilibrium solutions of a system like that in Eq. (4.3) (or, perhaps better (4.6)), i.e., the right-hand sides of them equal to zero, after the system has been put into its normal form. A $k$-parameter unfolding of (the normal form) $g$ (of the original problem) is a family $G(x,\lambda,\mu)$, for the $k$-parameter $\mu$ such that $G(x,\lambda,0)=g(x,\lambda)$. This is a perturbation of $g(x,\lambda)$ since,
$G(x,\lambda,\mu_{*})=g(x,\lambda)+p(x,\lambda),$ (4.10)
in the sense of qualitative equivalence. We arrive at the versal unfolding equation corresponding to the original problem, that is
$\dot{x}=G(x,\lambda,\mu).$ (4.11)
Setting $\mu=0$ in Eq. (4.11) we are back to Eq. (4.3) (or, (4.6)), that is to the original system written in the suitable dynamical system form that we motivated in subsection (4.1). The number $k$ of components of the auxiliary parameter $\mu$ is called the codimension of $g$.
Remark 4.2
We shall work with versal unfoldings since they also have a number of other benefits, including the determination of global (or quasi-global) properties using only local methods, as well as a novel way of thinking about the original problem as they impose a structure on the physical parameter space.
But perhaps the most remarkable property that is shared by the versal unfolding construction is that in distinction to the original system which it replaces, it is structurally stable, that is contains all stable perturbations of the original equation. In this sense the versal unfolding system represents in a somewhat exaggerated language, a (or perhaps ‘the’) ‘theory of everything’ of the original equation.
4.6 Bifurcation diagrams
A subtler property of the versal unfolding is related to the problem of the likelihood of the equations $g(x,\lambda)=0$, that is how to classify the qualitatively different types of these problems. One would naively expect an infinite number of them, one for each value of the unfolding vectorial parameter $\mu$ in the versal family $G(x,\lambda,\mu)$ (compare with our discussion of the central problem in Section 3).
However, the singularity theory notion of codimension provides an ingenious approach to this problem in that the likelihood of one such equation of a given qualitative type is higher the smaller the codimension of it. The bifurcation diagram, graph of the solution sets of the versal unfolding equation $G(x,\lambda,\mu)=0$, provides complete qualitative information about the original physical problem in this respect.
This is so because it describes not only the dynamics of the normal form of the gravity problem, $g(x,\lambda)=0$, but also that of all its perturbations through qualitative analysis of the dynamics of the versal unfolding equation.
In a sense, the normal form equation $g(x,\lambda)=0$ is the organizing centre of the dynamics of the original physical problem, since all possible bifurcation diagrams are generated by its versal unfolding.
4.7 Moduli
These are further types of parameters other than the distinguished and auxiliary (i.e., unfolding) ones. Moduli appear in the normal form and the versal unfolding equations as integer coefficients in front of some terms. The reason for their appearance is related to the existence of certain perturbations of $g$ in the versal unfolding which are topologically- but not $C^{\infty}$-equivalent to $g$ itself.
In this sense, one associates a topological codimension to the equilibrium of $g$ defined as its codimension minus the number of moduli coefficients.
The way to calculate the moduli is through algebraic relations that define the notion of genericity for versal unfoldings, these being conditions on the derivatives of $g$ of two kinds: nondegeneracy conditions involving derivatives with respect to the phase variables, and transversality conditions on derivatives with respect to the auxiliary parameters; the former conditions are need to ensure that the equilibrium of the system is not too degenerate, whereas the latter are required to assure that the parameters unfold the field in a generic way.
Although there are moduli coefficients in the constructions of most of the problems developed in the next Section, we do not use the topological notion of equivalence.
4.8 Symmetries
To define the notion of a symmetric dynamical system, let $\Gamma$ be a compact Lie group and consider its action of the vector space $V$ of the dependent variables $x$ of the problem $\dot{x}=g(x)$. We say that $g$ is $\Gamma$-equivariant if for all $\gamma\in\Gamma$, we have $g(\gamma x)=\gamma g(x)$. In particular, if $x_{0}$ is an equilibrium of the system, then so is $\gamma x_{0}$.
In this case, either $\gamma x_{0}\neq x_{0}$, and we have found a new equilibrium, or $\gamma x_{0}=x_{0}$ and $\gamma$ is a symmetry of the solution $x_{0}$ (this is useful because we may then enumerate only those equilibria not related by the symmetries of the problem).
There is a $\mathbb{Z}_{2}$-symmetry in some of the problems employed below, and indeed many other gravitational problems are equivariant with respect to some symmetry group.
For example, the Wianwright-Hsu equations for homogeneous cosmologies are $D_{3}$-equivariant where $D_{3}$ is the dihedral group, and this symmetry has important implications for the early evolution and transient behaviour of Bianchi dynamics, cf. [44].
In addition, the presence of symmetry in the problem of singularities treated in Section 5.3 has a very important implication for its codimension of that problem: since the linear part of the normal form is the zero matrix, it means that one needs to unfold using four parameters. However, the $\mathbb{Z}_{2}$-symmetry in that problem reduces the codimension of two so that we only need to unfold the field with only two parameters.
The use of equivariant singularity theory in these problems (cf. also below in the problem of the crease flow and the Friedmann-Lwmaître equations) plays a very decisive role in the progress that can be made given the very high complexity involved in the original equations.
4.9 Global bifurcations
Although such orbits have not been studied in gravity to any systematic degree, they indeed appear abundantly in some of the problems studied so far in gravitational bifurcation theory (for instance as saddle-connections).
Therefore global bifurcations constitute an inherent feature of gravitational phenomena that deserves to be further studied, in particular, the typical ways of how and when the breaking of such orbits occurs via homoclinic and heteroclinic bifurcations and new invariant sets are thereby created.
4.10 Physics of the versal unfoldings
The versal unfolding and its bifurcation diagrams have non-zero parameters everywhere, except at the origin of the parameter space $\mu=0$ which corresponds to the original system. All unfolding parameters are zero in the original system (and its normal form). Therefore the physics associated with versal families is expected to possess characteristics and properties foreign to those which one is accustomed with when working solely with the original system.
Below we assume that we have constructed a versal family corresponding to some original system (e.g., that found by some metric and matter source in relation to the Einstein equations) by starting from the latter and following the steps described in previous subsections.
Then new physics will emerge as a result of the cooperation of the following factors which are only linked to the versal unfolding.
  1. 1.
  2. 2.
    The versal unfolding system is structurally stable.
  3. 3.
  4. 4.
  5. 5.
    At a bifurcation set defined in parameter space, the system transfigures to a new stage of its evolution and becomes something else described by the phase diagram of the next stratum.
  6. 6.
We shall discuss various aspects of these effects and constructions in the next section. Also these results will put known properties of the original gravitation systems in a new context. We shall see that the physics of the versal unfolding is much more diverse than what one is accustomed with in the non-unfolded problem that one has started from.
The slightest deformation of the original system (now lying at the origin of the bifurcation diagram) will lead to a series of new and qualitatively inequivalent phase portraits and their continuous changes to other forms upon parameter variation.
As shown in great detail in the bifurcation diagrams of the next Section, that complexity is due to two factors: 1) the parameter space is higher-dimensional (rather than zero-dimensional that is a ‘point’ when the parameters are zero as in the original system), and 2) the centre manifold is also higher-dimensional (than zero).
More importantly, the Jordan form at the origin as is typical in gravitational problems is a degenerate matrix with two zero eigenvalues.
5 Gravitational bifurcations

### v2

@@ 3.1 The main problem: intuitive discussion
For a gravitating system modelled by a dispersive (i.e., structurally unstable) set of dynamical equations, there are two main, very different qualitatively, cases to consider:
@@
This is in sharp contrast to the possible behaviours of structurally stable, hyperbolic systems as discussed in the previous section of this paper. We say that gravitational equations are of infinite codimension in Case 1, and of finite codimension in Case 2. Although one does not rule out the existence of gravity problems belonging to Case 1, in any encountered problem of infinite-codimension one would typically obtain overdetermined problems and following R. Thom [56], chap. 3, one expects to be able to show that there exists a finite order such that ‘almost all’ enumerations at some smaller order become determinate at that order. This means that in this case higher-order terms beyond that order may be neglected, and we may treat this case as one of finite codimension [^5] There is also a more ‘pragmatic’ approach to versal unfoldings, where one truncates the jet at some order $k$ and considers perturbations with finitely many parameters for which the jet is $k$-determined only with respect to some given property (or properties). The unfolding will then be versal only with respect to those properties, not with respect to all. Although this is a very interesting alternative approach to versal unfoldings, it is not the one that is followed here. For more details on this approach to bifurcation theory (and for a number of more advanced/alternative treatments of topics of the more standard approach followed here), the reader is directed to Ref. [73]..
@@
The case of finite codimension, say equal to $s\leq 4$, is the main case we shall consider below. In fact, we shall advance a point of view and various results that point to the finite codimension answer, and for various reasons explained below-while we keep an open mind-we also expect that in most cases of interest the codimension for gravitational systems will be small (up to four).
Let us assume for concreteness that we are given a gravitating system governed by the Einstein equations (for example, a Friedmann-Robertson-Walker (FRW) universe with a perfect fluid source-the following argument remains applicable in any modified gravity and effective string theory situation, etc). There are two basic issues on which our approach to gravitational bifurcations focuses: firstly, the question about finite determinacy: to what extent the low-order terms in a Taylor expansion of the gravitational bifurcation problem completely determine its qualitative behaviour regardless of the higher-order terms that may be present? Secondly, the question of unfolding: how does a gravitational bifurcation depend on parameters? The natural area to formulate both questions and search for complete answers is provided by the elaborate mathematical framework of bifurcation theory. Armed with the ideas of the previous section, we imagine that when an equilibrium $(x,\mu)=(0,0)$ at the origin of a parametrized family of systems undergoes a bifurcation at $\mu=0$, then the flow for $\mu$ near zero and $x$ near zero is qualitatively different to that near $x=0$ at $\mu=0$.
We may now describe the main objectives of the present approach in a somewhat more precise language. In the process we shall discover how the main methods of bifurcation theory discussed in Section 2 fit into the general problem of gravitation.
3.1.1 Objectives
In any gravitational context for example, cosmology, black holes, modified gravity, or effective string theory, we aim at a complete realization of the program of bifurcation theory for a variety of gravitational problems. This program is one of deciding the following goals:
  1. 1.
    Determine the nature and types of bifurcational behaviour and decide on the degree of degeneracy of the problem.
  2. 2.
    Resolve the problem of higher-order terms and the issue of parameters.
  3. 3.
    Discover the physical implications of the constructed versal families associated with the problem.
This program will be accomplished below by achieving progress in three directions that together offer a complete portrait of all possible qualitatively different behaviours for a given gravitational problem (the question of how to accomplish these objectives was analysed in Section 2, while examples of how these objectives have worked out so far will be given below.
First objective: For typical gravitation problems, demonstrate that they are bifurcation theory problems.
This objective deals with the first main difficulty to realize the bifurcation theory program, namely, to identify the type of degenerate behaviour present in the given gravity problem and separate it from any remaining hyperbolic one. We have to demonstrate that the given set of gravity equations indeed constitute a bifurcation problem. In this regard, we need to solve the following problems:
@@ 3.1.1 Objectives
Second objective: Given the results of Objective-1 for a given gravity problem, provide complete answers to the finite determinacy and unfolding questions.
@@
Third objective: Develop the physics of versal unfoldings associated with the original gravity problem.
While the first two objectives are almost purely mathematical and lead to the constructions of the topological normal forms and the bifurcation diagrams, in this objective the main question is to discover the physical implications stemming from these constructions as regards the original physical problem. The importance of this objective naturally leads to new physical effects totally absent in the original system. This clearly follows from the fact that the resulting constructions from the first two objectives are very closely interlocked with, and follow from, the degeneracies of the original system, all of which are by necessity ignored in a hyperbolic treatment. To motivate this objective, let us start by emphasizing the fact that there are three kinds of parameters in the versal unfolding (and associated bifurcation diagrams):
@@ 3.1.2 Importance
3.1.3 Gravitational selection and completion mechanisms
It follows that a bifurcation theory approach to gravitation leads to the discovery of novel aspects of gravitation that cannot be found without it, and also provides the means for their further study. It implies possibly observable effects and signatures, definite predictions which would be difficult to capture otherwise but ready to be confirmed or falsified. This aspect of the present work will also be important in physical cosmology starting from bifurcation results. This is so for the following three main reasons:
@@ 3.1.3 Gravitational selection and completion mechanisms
  • Novel mathematical problems appear by necessity. Bifurcation theory contains a purely algebraic approach to gravitational problems in which the resulting equations fully determine the solutions. For example:
@@
    Armed with the versal unfolding results, the question now arises as to how the novel branches - and not just the solutions sitting at the origin of the parameter space - are lifted, and by which mechanism are selected, by the ‘mother’ partial differential equation (for example the Einstein equation, which was ‘collapsed’ to the ‘original problem’ due to a symmetry reduction process as discussed earlier). This also includes the general question of stability (in the PDE sense) and extends the problem to that of determining the gravitational selection mechanism which becomes responsible for eventually choosing which branch to keep and which to leave in the process.
@@
This is where the importance of a gravitational bifurcation approach lies. It represents the first few steps towards a more comprehensive approach of ‘gravitational selection and completion’ by generic perturbations, and independently, an approach to the problem posed above. For it provides a reliable path from the large body of known and well-understood results related to a hyperbolic approach, to the description of a consistent (i.e., structurally stable) framework, a new but yet unknown structure sought-for in this approach to general relativity, which contains all ‘surviving’ versal unfoldings of a given original idealization. For example, what are the versal unfoldings and the precise relations between them for the Einstein-Maxwell, Einstein-Dirac, or Wheeler-DeWitt equations?